OpenStudy (anonymous):

Please help! Will fan and medal! eliminate the parameter x = t + 4, y = t^2

OpenStudy (anonymous):

Ok. In first equation, leave t alone: \[x = t + 4\] \[t = x - 4\] Afther this, put that 't' in the second equation. \[y = (x - 4)^2\] Thats it.

OpenStudy (anonymous):

I believe it's y = x^2 - 8x + 16. The answer you gave isnt one of my options =\ My options are: y = x^2 - 8x + 16 y = x^2 + 16 y = sqrt(x-4) y = sqrt x + x + 4

OpenStudy (anonymous):

yes, you just needed to arrange this expression. Your answer is right.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

anytime